OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3.

OpenStudy (anonymous):

Point-slope form looks like y-y1=m(x-x1) x1=-1 y1=6 m=-3 So y-6=-3(x+1)